OpenStudy (anonymous):

prove the identity: cos((pi/2)-x)= sinx

OpenStudy (anonymous):

By cos(α-β)=cosα·cosβ+sinα·sinβ, cos(pi/2-x)=cos(pi/2)·cos(-x)+sin(pi/2)·sin(-x) =0·cos(-x)+(-1)·sin(-x) =0+(-1)·sinx =-sinx